UT 20 November 2018

OSA: J. Williams

ISA: B. Rothberg

ISp: G. Bechetti

Mtn Manager: M. Wagner

AO Support (ARGOS): S. Walsh

AO Support (FLAO): D. Miller

Laser Dude: G. Rahmer

Software Support: A. Colson, I. Sola

luciN.20181120.XXXX

UT 06:46 - LUCI GUIs started up - init all ran

L1 0001-0002 Test Darks

L2 0015-0016 Test Darks

UT 07:04 - reconfiguring for Bino LUCI - starting up WFS/ARGOS

UT 07:09 - pointing check

UT 07:26 - collimation script sent - WFS shows 0.5"

UT 07:29 - loops closed DX - tip tilt only - FLAO way out of focus

L1 0003 - test exposure - 0.45" in center of field (SL)

UT 07:36 - ripples seen in secondary on the FLAO GUIs - Doug is taking telemetry data from the FLAO side

L2 0017 - test L2 exposure to see state of system - same elongation seen NE to SW 2.51x13

L2 0018- elongation still the same - the last LGS command sent has been frozen into the mirror (when gains were non-zero) 2.51x13

L2 0019 - loop open - doing tip/tilt only - less elongated - more rounder - Doug M. took telemetry data (2.51x5)

UT 07:45 - trying to sharpen laser spots

L2 0020 - taking another exposure (2.51x5) - lots of coma

- coma likely due to hexapod having moved at some point

UT 07:53 - sending collimation scripts again to start from scratch with testing

UT 07:57 - collimated again

L2 0021 - active optics collimated image (2.51 x5)

UT 08:00 - sending preset for AROGS DX / SL SX - configuring only tip tilt on DX side

UT 08:01 - pointing - guidestar on SX missed in the acquisition

- at elevation 88

UT 08:13 - transiting - so in a holding pattern until the rotators slow down

UT 08:15 - sent acitve optics collimation preset

UT 08:23 - pointing check nearby M-star

UT 08:25 - mountain connection dropped out

UT 08:29 - sending collimation preset

UT 08:30 - now missed both stars

UT 08:32 - pointing check again

UT 08:36 - resending active collimation script again to Pal2

UT 08:40 - L1 0004 - test exposure SL (2.51x13)

L2 0022 - test exposure L2 (2.51x5)

UT 08:43 - sending ARGOS science script

-lost communication with slope computer

UT 08:45 -resending preset

UT 08:46 -lost mountain communication again

L2 0023 - loops closed tip-tilt only - image shows more round, but defintiely triangle

UT 08:49 - lasers jumping around

L1 0005 - test exposure to compare SL active optics - elongation on L1 images

L1 0006 - another LUCI1 test image - looks a little better

UT 08:52 -Doug M. notes pyramid image is way out of focus - could be doing a weird tip-tilt causing a problem (CCD 39)

-changed PSGUI right z-global offset ~0.3

L2 0024 - test exposure after offset - some improvement in image quality

-next step +0.1 low order gain

L2 0025 - slight elongation - looks good though

L2 0026 -low order +0.2

L2 0027 - tiptilt 0.75

L1 0007 - comparison SL image

L1 0008 - comparison SL image

L2 0028 high order added 0.1 gain - definitely see strong elongation

L2 0029 -sending gain back to 0 for high order

L1 0009, L2 0029 - L2 image shows the same thing imprinted even though high order turned back down

open loops will restore the flat

UT 09:08 - tip-tilt 1.0, low order 0.2, 0 high order L2 0030 - test exposure - definitely elongated diagonally again

UT 09:16 - bumping high order 0.1 - L2 0031 ( comparions L1 0010)

UT 09:20 - opening loops

-thin shell is creating spikes

UT 09:30 - aligning FM4 on LUCI-1 (L1 0011-0013)

UT 09:34 - aligning FM4 on LUCI-2 (L2 0032-0033)

UT 09:37 - resending preset, lost star on the SX and AORef on the right

UT 09:38 - pointing check

UT 09:45 - sending collimation script for Palomar 2

UT 09:48 - sending preset to Palomar 2 (SL/ARGOS)

UT 09:54 - testing the ripple effect and if it can be removed

open loop caused rip, recovery didn't work, had to cycle IDL and fast diagnostic process - powering on system and will set flat (happens occassionally - 2-3x/night with LBTI)

UT 10:00 - resending preset

L2 0034 - test image - stars look like sausages

UT 10:22 - moving to HCG 31 as a test of low elevation

- pointing check

UT 10:25 - sending HCG31 collimation script

UT 10:29 -sending preset SL/ARGOS to HCG 31

Acuator Failure in AOS GUI

-satellite avoidance thinks we are still slewing, we are not

UT 10:32 - resending preset

- had rolled back software to version before elevations < 45 were allowed

-right side failed again with ARGOS arbitrator issue

UT 10:36 - went back to collimation preset - due to erroneous communication from Laser Dude

UT 10:39 - laser dude firing at clouds

UT 10:44 - stopping ARGOS, switching to ESM

UT 10:49 - sending preset

UT 10:52 - didn't see guidestars or AORef star (R=12, 10.9, respectively)

-could be due to pointing and reverting back to IE/CA older values

UT 11:00 - resending preset - can't find pointing star either

UT 11:04 - going to try NGC 2782

UT 11:08 - can see something for pointing stars (for now)

UT 11:10 - another drop out from mountain

UT 11:12 - sending SL/ESM Science preset to NGC 2782

UT 11:14 - timeout error on FW2 - AO error

UT 11:16 - resending preset to NGC 2782 SL/ESM

UT 11:18 - AO supervisor problem on DX, stopping and restarting AO

UT11:24 - resending preset SL/ESM to NGC 2782

UT 11:28 - lost last vestages of photons on left too, put in blind filters, will sit hoping it gets better (super thick clouds)

UT 11:39 - resending preset

UT 11:46 - running ESM - clouds cleared overhead for now

optimized gain timed out on IIF

RunAO failled on offset

-attempting to trick it into running

UT 11:54 - resending preset again SL/ESM - stop AO didn't get sent - caused DX RIP

UT 12:03 - preset failed - clouds thickining

UT 12:20 - M5 star bright in guider, trying preset again (NGC 2782 SL/ESM)

UT 12:22 - no stars appeared in the acquisition

UT 12:27 -nearby pointing star - OK - resending NGC 2782

L1 0014 -

L2 0035-

UT 12:41 - returned to field but couldn't find AO Ref Star

-failed to see anything back in the field

DX AdSec being restarted

UT 12:46 - Think we are done - clouds thick - -12 twilight is 10min away

No useful (ESM, ARGOS, SL comparison) data taken tonight

-- BarryRothberg - 20 Nov 2018
Topic revision: r1 - 20 Nov 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback